OpenStudy (anonymous):

Data downloaded from a permanent count station yielded that the evening peak hour traffic averaged 1915 (vph) and had a standard deviation of 430 (vph).  The data appear to follow a normal distribution. Previous analyses found that the roadway has a capacity of 2275 (vph).       a.) What is the probability that the hourly traffic will exceed capacity?  

OpenStudy (shaik0124):

since the data is normal distribution median is 1915 ( avg is equal to median when data is equally spaced)

OpenStudy (shaik0124):

in normal distribution mean ,median and mode all are equal so the next hour traffic will be 1915+430 =2345

OpenStudy (shaik0124):

consider this set s={2,3,4,5} difference beween each number is 1 i.e,equally spaced now mean =(2+3+4+5)/4 =3.5 meadian =3+4)/2 =3.5

OpenStudy (shaik0124):

dou have answer choices to this question

OpenStudy (anonymous):

No, but I think I got the answer.

OpenStudy (shaik0124):

so how did u solve

OpenStudy (anonymous):

 So we want to probability that x > 2275. P(x>2275) z = x-u/o = 2275-1915/430 = 0.84 So now we have P(z>0.84) = 1-P(z<0.84) = 1- 0.7995 We get 0.7995 from something called the z score table. Here's a video with examples. https://www.youtube.com/watch?v=mai23vW8uFM